Blocks A (mass 5.00 kg ) and B (mass 14.50 kg ) move on a frictionless, horizontal surface. Initially, block B is at rest and block A is moving toward it at 9.00 m/s . The blocks are equipped with ideal spring bumpers. The collision is head-on, so all motion before and after the collision is along a straight line. Let +x be the direction of the initial motion ofA.

Part A

Find the maximum energy stored in the spring bumpers and the velocity of each block at that time.
Find the maximum energy.

Ub___ J

Part B

Find the velocity of A

Part C

Find the velocity of B

Part D

Find the velocity of each block after they have moved apart.
Find the velocity of A

vA___m/s

Part E

Find the velocity of B

vB___m/s

some one did it for me like this

initialy the bumpers have maximum energy stores

then maximum energy stored is U = 0.5*5*9^2 = 202.5 J

Part B

VA = 9 m/s

Part C

VB = 0 m/s


Part D

VA = (m1-m2)*u1/(m1+m2)= (5-14.5)*9/(5+14.5) = -4.4 m/s

PartE

V2 = u1+v1 = 9-4.4 = 4.6 m/s

but its incorrect... can anyone help me pleaseeee
Answer #1

First we have to find velocity of its centre of mass

V_{cm } = \frac{m_1v_1 + m_2 v_2}{(m_1+ m_2)}

V_{cm } = \frac{5*9 + 14.5*0}{(5+ 14.5)} = 2.31 m/s

Now Kinetic energy of Centre of mass of the system

KE = \frac{1}{2}(m_1+m_2)V_{cm}^2

KE = \frac{1}{2}(5 + 14.5)(2.31)^2 = 51.92 J

Initial total kinetic energy of two blocks

KE_i = \frac{1}{2}m_1v_1^2 + \frac{1}{2}m_2v_2^2

KE_i = \frac{1}{2}5*9^2 + \frac{1}{2}14.5*0^2 = 202.5 J

Now the maximum energy stored in spring is given by

KE_{max} = 202.5 - 51.92 = 150.58 J

PART B)

At the same time when energy stored in the spring is maximum then both the blocks will move with common velocity

V_a = V_{cm} = 2.31 m/s

PART C)

Similarly as stated above

V_b = V_{cm} = 2.31 m/s

PART D)

when two blocks are moved apart

V_{af} = \frac{(m_1 - m_2)u_1}{m_1+m_2}

V_{af} = \frac{(5 - 14.5)*9}{5 +14.5} = -4.38 m/s

PART E)

V_{bf} = \frac{2m_1v_1}{m_1+m_2}

V_{bf} = \frac{2*5*9}{5+14.5} = 4.62 m/s